OpenStudy (anonymous):

(y3 - 125) ÷ (y - 5)

OpenStudy (anonymous):

(y³ - 125) / (y - 5) Use (a³- b³) identity i.e a^ + ab + b^ [ ^ = 2] = (y³ - 5³ ) / (y - 5) = (y - 5) ( y^ + 5y + 5^) / (y - 5) [^ = 2] = y^ + 5y + 25